Agyenim is concerned about his exposure because he believes that there are two possible scenarios: (1) the pound will depreciate by 3 percent over the next month or (2) the pound will appreciate by 2 percent over the next month. There is a 70 percent chance that Scenario 1 will occur. There is a 30 percent chance that Scenario 2 will occur.

Agyenim notices that the prevailing spot rate of the pound is GHS 8.1 and the one- month forward rate is GHS 8.6. Agyenim can purchase a put option over the counter from a securities firm that has an exercise (strike) price of GHS 8.6, a premium of GHS0.025, and an expiration date of one month from now. Determine the amount of cedis received by the Best Pineapple Company under each of the two exchange rate scenarios if:

a) The receivables to be received in one month are not hedged.

b) A put option is used to hedge the receivables in one month.

c) A forward hedge is used to hedge the receivables in one month.



1
Expert's answer
2021-08-10T11:45:17-0400



a)in this case, just the debtor will simply depend on the exchange rate of the pound sterling:

"20 000-20000\times0.03=19,400"

"19 400\times8.1=157140"

"19400\times8.6=166 840"


"20 000\times0.02+20 000=20 400"

"20 400\times8.1=165 240"

"20 400\times8.6=175 440"

b)

A put option is a contract for the right to sell an asset before a certain date in the future at a price and quantity determined at the current moment.

(1)"8.1-8.1\times0.03=7.857"

"7.857\times20 000=157140"

(2)"8.1+8.1\times0.2=8.262"

"8.262\times20 000=165 240"


c)A forward contract is not a standardized contract for the delivery of the underlying asset in the future at a predetermined price, not an exchange instrument.

(1)"8.6-8.1\times0.03=8.357"

"8.357\times20 000=167 140"

(2)"8.6+8.1\times0.2=8.762"

"8.762\times20 000=175 240"